OpenStudy (anonymous):

solve for x 2x-3/3 +5-x/2 =3

hero (hero):

\[\frac{2x-3}{3} + \frac{5 - x}{2} = 3\]

hero (hero):

Is that it?

OpenStudy (anonymous):

yes

hero (hero):

\[\frac{2x - 3}{3} + \frac{5 - x}{2} = \frac{6}{2}\]\[\frac{2x - 3}{3} = \frac{6}{2} - \frac{5 - x}{2}\]\[\frac{2x - 3}{3} = \frac{6 - 5 + x}{2} \]\[\frac{2x - 3}{3} = \frac{1+ x}{2}\]\[2(2x-3) = 3(1+x)\] can you finish the rest?

OpenStudy (anonymous):

i dont understand

hero (hero):

6/2 = 3, right?

OpenStudy (anonymous):

yes

hero (hero):

Okay, so I replaced 3 with 6/2. Now \(\frac{5 - x}{2}\) and \(\frac{6}{2}\) have the same denominator.

hero (hero):

So I subtract \(\frac{5-x}{2}\) from both sides. Once I do that, I have to subtract: \(\frac{6}{2} - \frac{5 -x}{2}\)

hero (hero):

When I do that, I end up with \(\frac{1 + x}{2}\)

hero (hero):

So now what's left is: \(\frac{2x-3}{3} = \frac{1+x}{2}\) Which you use cross multiply to get \(2(2x-3) = 3(1+x)\)

hero (hero):

Can you finish solving from there?

OpenStudy (anonymous):

not sure but I'll try

OpenStudy (anonymous):

thanks

hero (hero):

Did you solve for x? If so, what did you get?

OpenStudy (anonymous):

is it 1

hero (hero):

No